# Calibrating dimension reduction hyperparameters in the presence of noise

### 05 Dec 2023 v1

@@ 6 Discussion
We have illustrated the importance of acknowledging noise when performing dimension reduction by studying the roles perplexity and n_neighbors play in overfitting data. When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended. When using just the signal, however, our experiments indicated that larger perplexities performed better. Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability. These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics (Chu et al., 2022).
@@ 7 Data Availability
All data and example code are freely available at https://github.com/JustinMLin/DR-Framework/.

### 30 Jan 2024 v3

@@ 6 Discussion
We have illustrated the importance of acknowledging noise when performing dimension reduction by studying the roles perplexity and n_neighbors play in overfitting data. When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended. When using just the signal, however, our experiments indicated that larger perplexities performed better. Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability. These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics [5].
@@ 7 Data Availability
All data and code are freely available at https://github.com/JustinMLin/DR-Framework/.
Supplementary Information
SI Simulated Examples
SI.1 Trefoil Plots
Figure S1: Trefoil Plots
SI.2 Mammoth Plots
Figure S2: Mammoth Plots
SII Practical Examples
SII.1 UMAP Plots (CyTOF)
Figure S3: UMAP Plots
SII.2 scRNA Dataset
This is a dataset of induced pluripotent stem cells generated from three different individuals [16]. The original data includes 864 units and 19,027 readings per unit. To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied. This reduced the dimension to 5,431. A PCA pre-processing stop further reduced the dimension to 500, which still retained 88% of the variance of the log-transformed data. The signal was first taken to be the first five principal components, then the first 10 principal components. Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension. This is due to the inherent randomness of the t-SNE algorithm.
Figure S4: scRNA Plots (r = 5)
Figure S5: scRNA Plots (r = 10)
SII.3 Microbiome Dataset
[2] compares the faecal microbial communities from 22 subjects using complete shotgun DNA sequencing. The original data contained 280 samples and 553 genera. To deal with a large number of near-zero readings, columns containing a large proportion of values less than $10^{-6}$ (60% or more) were removed. This reduced the dimension to 66. A PCA pre-processing was used to center and re-scale the data. The signal was first taken to be the first five principal components, then the first eight principal components. Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension. This is due to the inherent randomness of the t-SNE algorithm.
Figure S6: Microbiome Plots (r = 5)
Figure S7: Microbiome Plots (r = 8)